OpenStudy (anonymous):

hi, how do you work out 1/k! by using the ratio test?

OpenStudy (anonymous):

the ratio test uses {1/(k+1)!}/(1/k!)=1/(k+1) and takes the limit as k tends to infinity lim(1/k+1)=0 (by the Archimedes rule)

OpenStudy (anonymous):

hope it is clear